# Alignment-Sensitive Minimax Rates for Spectral Algorithms with Learned Kernels

### v1

@@ 8.3 Connection to Random Design Linear Regression
The treatment of analysis developed in this section covers an important model, the random-design linear regression in contrast to the fixed-design linear regression. This is because linear regression can be viewed as a RKHS regression w.r.t. any positive definite linear kernels $\mathbf{k}(x,x^{\prime})=x^{\top}\mathbf{K}x^{\prime}$ for $x,x^{\prime}\in\mathbb{R}^{p}$, where $\mathbf{K}\in\mathbb{R}^{p\times p}$ is positive definite.
@@
where $\mathbf{V}$ is the matrix with columns formed by $v_{j}$ and $\mathbf{\Lambda}=\operatorname{Diag}(\ldots\lambda_{i}\ldots)$.
@@
Hence, $(\{\lambda_{i}\},\{\psi_{j}\})$ is the eigen-pair for $\mathbf{k}$.
@@ 8.4 Numerical Illustration
with $D=80$. As $\alpha$ increases, the leading $D$ eigenvalues become progressively magnified, with the largest index having the most significant increase. Consequently, the modified kernel will emphasize more on directions that received less energy of the signal and the optimal KPCPE should require selecting more principal components.
@@ 9 Measuring Alignment via ESD
We illustrate how the notion of ESD can be used to measure the alignment between the signal and the kernel (spectrum).
9.1 An example of comparing signal-spectrum alignments
The following is a simple example to illustrate the idea of comparing the signal-spectrum alignments for different spectra discussed in Section˜3.2.
@@ 9.1 An example of comparing signal-spectrum alignments
(1) The first $k$ largest eigenvalues of $\bm{\lambda}^{(1)}$ are located on $S$;
(2) The first $k$ largest eigenvalues of $\bm{\lambda}^{(2)}$ are located on $S^{c}=[d]\setminus S$.
Intuitively, $\bm{\lambda}^{(1)}$ aligns better with $\bm{\theta}^{*}$ than $\bm{\lambda}^{(2)}$. However, a quantitative analysis is not obvious without using the notion of ESD.
Firstly, we note that the effective dimensions [51] for these two spectra are the same because the sets of eigenvalues are the same. Similarly, the covariance splitting indices $k^{*}$ [5] are the same for the two spectra. We confirm that signal-agnostic complexity measures do not distinguish the signal-spectrum alignments for these two spectra.
Next, we consider the ESD and span profile. Rigorously, we can show for any $\tau$,
@@
$r(\tau)={\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(1)}}(\tau)}/{\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(2)}}(\tau)}\leq s/(d-s)\ll 1$
In view of Theorem˜3.3, this suggests that the minimax estimation using $\bm{\lambda}^{(1)}$ is substantially than using $\bm{\lambda}^{(2)}$ when the noise level is small. Therefore, spectral estimators using $\bm{\lambda}^{(1)}$ is preferred.
@@ 9.2 Pathwise ESD for Learned Kernels
Let $\mathbf{k}_{t}$ be the kernel learned at training time $t$, with eigenvalues $\{\lambda_{j}(t)\}$ (sorted decreasing) and $L^{2}(\mathcal{X},\mu)$-orthonormal eigenfunctions $\{\psi_{j}^{(t)}\}$. To understand how the signal-kernel alignment evolves, we define the pathwise ESD as
@@
Let $\mathbf{H}_{t}(k):=\frac{1}{k}\sum_{i>k}\left[\theta^{*,(t)}_{i}\right]^{2}$. If training aligns the leading eigenfunctions $\psi_{j}^{(t)}$ better with $f^{*}$, then $\{\theta_{j}^{*,(t)}\}$ concentrate more on leading indices, and thus $\mathbf{H}_{t}(k)$ decreases for all $k$, which implies the decrease in $d^{\dagger}(t)$.
@@
Each covariate coordinate is drawn independently from $\{\pm 1\}$, so $\bm{\Sigma}_{x}=\mathbb{E}(XX^{\top})=\mathbf{I}_{p}$ and $\|X\|_{\infty}=1$. We set $p=900$, and the true parameters are set as follows: $\beta^{*}$ follows a power-law decay with $\beta^{*}_{j}=j^{-1.1}$ for $1\leq j\leq 200$ and $\beta^{*}_{j}=0$ for $j>200$. The response is $Y=\langle\beta^{*},X\rangle+\varepsilon$ with $\varepsilon\sim N(0,\sigma_{0}^{2})$ and $\sigma_{0}=0.1$.
We draw $n=1000$ samples and train a deep linear network with $D=4$ hidden affine maps without bias using full-batch Adam with learning rate $10^{-4}$. The hidden weight matrices of the network are $\mathbf{W}_{\ell}(t)\in\mathbb{R}^{p\times p}$ for $\ell=1,\ldots,D$ (using a near-identity initialization) and the weight of a final linear layer is $w(t)\in\mathbb{R}^{p}$.
The estimated function at time $t$ is given by $f_{t}(x)=w(t)^{\top}\mathbf{A}(t)x$, where $\mathbf{A}(t):=\mathbf{W}_{D}(t)\cdots\mathbf{W}_{1}(t)$. We form the learned kernel $\mathbf{k}_{t}(x,x^{\prime})=\langle\mathbf{A}(t)x,\mathbf{A}(t)x^{\prime}\rangle=x^{\top}\mathbf{G}_{t}x^{\prime}$ where $G_{t}:=\mathbf{A}(t)_{t}^{\top}\mathbf{A}(t)_{t}$. We then follow the derivation in Section˜8.3 and define the ESD $d^{\dagger}(t)$ of $f^{*}$ w.r.t. the kernel $\mathbf{k}_{t}$. Since $\|X\|_{\infty}=1$ $\mu$-a.s., we have $\|f\|_{\infty}^{2}=\|\beta\|_{1}^{2}=\|\beta\|_{1}^{2}$; this is used in computing the effective noise level.
@@
Figure˜5 shows that adaptive representation learning progressively reduces the ESD $d^{\dagger}(t)$ in $t$ along with the true risk. This confirms that ESD captures the evolving alignment between signal and kernel.
@@ C.5 Auxiliary Lemma
Appendix D Related works on principal component regression
As discussed in Section˜3.1, the PC estimator serves as a motivating example for the concepts of ESD and span profile due to its clear illustration of bias-variance trade-offs. However, the ESD and span profile are designed to characterize the intrinsic difficulty of generalization arising from signal-kernel alignment, and their definition do not rely on the PC estimator. Nonetheless, the analysis of PC estimators, particularly in high-dimensional linear regression, been an active area of recent research. Below, we briefly summarize some relevant contributions to provide context.
D.1 Proportional Asymptotic Limits
Several studies analyze Principal Component Regression (PCR) in the proportional asymptotic setting where the dimension $p$ and sample size $n$ grow with $p/n\!\to\!\gamma$. In this regime, Xu and Hsu [46] studied the limiting risk of PCR with Gaussian designs with diagonal covariance. They assume polynomially decaying eigenvalues or convergent empirical spectra and an isotropic prior and reveal a “double-descent” risk curve. In a related vein, Wu and Xu [45] extend the analysis to general covariance $\Sigma_{x}$ and an anisotropic prior satisfying $\mathbb{E}\beta_{*}\beta_{*}^{\!\top}\!=\!\Sigma_{\beta}$. They also derive an exact risk expression and demonstrate how “misalignment” between $\Sigma_{x}$ and $\Sigma_{\beta}$ affects risk; here “alignment” refers to concordance between the orderings of their eigenvalues. Both studies assume knowledge of the eigenvectors of the population covariance matrix $\Sigma_{x}$ to construct the oracle PCR. Gedon et al. [17] analyze the limiting risk of PCR under a latent factor model and explore the effect of distribution shift. Green and Romanov [20] derive the exact limits of estimation risk, in-sample prediction risk, and out-of-sample prediction risk of PCR under the assumption that both the empirical distribution of the spectrum and the distribution of mass of the true signal over the eigenspace of $\Sigma_{x}$ converges weakly.
@@ D.2 Non-Asymptotic Analysis
Complementary research develops non-asymptotic guarantees. Agarwal et al. [1] derive finite-sample upper bounds on prediction error using $\|\beta^{*}\|_{1}^{2}$ and the rank of design matrix under latent factor models and explore the robustness of PCR to noise and missing values in the observed covariates. Bing et al. [9] consider PCR with adaptively selected number of components under latent factor models and provide alternative finite-sample risk bounds using $\|\beta^{*}\|_{2}^{2}$. Huang et al. [21] derived non-asymptotic risk bounds for PCR in more general settings by analyzing the alignment between population and empirical principal components. Hucker and Wahl [22] derive non-asymptotic error bounds for PCR in kernel regression.

### v2

@@ 8.3 Connection to Random Design Linear Regression
The analysis in this section also covers an important case: random-design linear regression (as opposed to fixed-design linear regression). This is because linear regression can be viewed as an RKHS regression w.r.t. any positive-definite linear kernels $\mathbf{k}(x,x^{\prime})=x^{\top}\mathbf{K}x^{\prime}$ for $x,x^{\prime}\in\mathbb{R}^{p}$, where $\mathbf{K}\in\mathbb{R}^{p\times p}$ is positive definite.
@@
where $\mathbf{V}$ is the matrix with columns formed by $v_{j}$ and $\mathbf{\Lambda}=\operatorname{Diag}(\lambda_{1},\dots,\lambda_{p})$.
@@
Hence, $(\{\lambda_{i}\},\{\psi_{j}\})$ is the eigenpair for $\mathbf{k}$.
@@ 8.4 Numerical Illustration
with $D=80$. As $\alpha$ increases, the leading $D$ eigenvalues become progressively magnified, with the largest index having the most significant increase. Consequently, the modified kernel places more emphasis on directions that receive less signal energy, so the optimal KPCPE selects more principal components.
@@ 9 Measuring Alignment via ESD
We illustrate how the notion of ESD can be used to measure the alignment between the signal and the kernel.
9.1 An example comparing signal-spectrum alignment
The following simple example illustrates how to compare signal-spectrum alignment across the spectra discussed in Section˜3.2.
@@ 9.1 An example comparing signal-spectrum alignment
(1) The $k$ largest eigenvalues of $\bm{\lambda}^{(1)}$ are located on $S$;
(2) The $k$ largest eigenvalues of $\bm{\lambda}^{(2)}$ are located on $S^{c}=[d]\setminus S$.
Intuitively, $\bm{\lambda}^{(1)}$ aligns better with $\bm{\theta}^{*}$ than $\bm{\lambda}^{(2)}$. However, a quantitative analysis is not straightforward without using the notion of ESD.
First, we note that the effective dimensions [52] is the same for both spectra because they share the same sets of eigenvalues. Similarly, the covariance-splitting index $k^{*}$ [5] is the same for both spectra. Thus, these signal-agnostic complexity measures do not distinguish signal-spectrum alignment between the two spectra.
Next, we consider the ESD and the span profile. Rigorously, we can show for any $\tau$,
@@
$r(\tau)={\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(1)}}(\tau)}/{\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(2)}}(\tau)}\leq s/(d-s)\ll 1.$
In view of Theorem˜3.3, this suggests that the minimax risk under $\bm{\lambda}^{(1)}$ is substantially lower than under $\bm{\lambda}^{(2)}$ when the noise level is small. Therefore, spectral estimators using $\bm{\lambda}^{(1)}$ are preferred.
@@ 9.2 Pathwise ESD for Learned Kernels
Let $\mathbf{k}_{t}$ be the kernel learned at training time $t$, with eigenvalues $\{\lambda_{j}(t)\}$ (sorted in decreasing order) and eigenfunctions $\{\psi_{j}^{(t)}\}$ that are orthonormal in $L^{2}(\mathcal{X},\mu)$. To understand how the signal-kernel alignment evolves, we define the pathwise ESD as
@@
Let $\mathbf{H}_{t}(k):=\frac{1}{k}\sum_{i>k}\left[\theta^{*,(t)}_{i}\right]^{2}$. If training aligns the leading eigenfunctions $\psi_{j}^{(t)}$ better with $f^{*}$, then $\{\theta_{j}^{*,(t)}\}$ concentrate more on the leading indices, and thus $\mathbf{H}_{t}(k)$ decreases for all $k$, which implies a decrease in $d^{\dagger}(t)$.
@@
Each covariate coordinate is drawn independently from $\{\pm 1\}$, so $\bm{\Sigma}_{x}=\mathbb{E}(XX^{\top})=\mathbf{I}_{p}$ and $\|X\|_{\infty}=1$. We set $p=900$, and specify the true parameters as follows: $\beta^{*}$ follows a power-law decay with $\beta^{*}_{j}=j^{-1.1}$ for $1\leq j\leq 200$ and $\beta^{*}_{j}=0$ for $j>200$. The response is $Y=\langle\beta^{*},X\rangle+\varepsilon$ with $\varepsilon\sim N(0,\sigma_{0}^{2})$ and $\sigma_{0}=0.1$.
We draw $n=1000$ samples and train a deep linear network with $D=4$ hidden affine layers without bias using full-batch Adam with learning rate $10^{-4}$. The hidden weight matrices of the network are $\mathbf{W}_{\ell}(t)\in\mathbb{R}^{p\times p}$ for $\ell=1,\ldots,D$ (using a near-identity initialization), and the weight of the final linear layer is $w(t)\in\mathbb{R}^{p}$.
The estimated function at time $t$ is given by $f_{t}(x)=w(t)^{\top}\mathbf{A}(t)x$, where $\mathbf{A}(t):=\mathbf{W}_{D}(t)\cdots\mathbf{W}_{1}(t)$. We form the learned kernel $\mathbf{k}_{t}(x,x^{\prime})=\langle\mathbf{A}(t)x,\mathbf{A}(t)x^{\prime}\rangle=x^{\top}\mathbf{G}_{t}x^{\prime}$, where $G_{t}:=\mathbf{A}(t)^{\top}\mathbf{A}(t)$. We then follow the derivation in Section˜8.3 and define the ESD $d^{\dagger}(t)$ of $f^{*}$ w.r.t. the kernel $\mathbf{k}_{t}$. Since $\|X\|_{\infty}=1$ $\mu$-a.s., we have $\|f^{*}\|_{\infty}^{2}=\|\beta^{*}\|_{1}^{2}$; this is used in computing the effective noise level.
@@
Figure˜5 shows that adaptive representation learning progressively reduces the ESD $d^{\dagger}(t)$ over training time, along with the true risk. This confirms that ESD captures the evolving alignment between signal and kernel.
@@ C.5 Auxiliary Lemma
Appendix D Related Work on Principal Component Regression
As discussed in Section˜3.1, the PC estimator serves as a motivating example for the concepts of ESD and span profile due to its clear illustration of bias-variance trade-offs. However, the ESD and span profile are designed to characterize the intrinsic difficulty of generalization arising from signal-kernel alignment, and their definitions do not rely on the PC estimator. Nonetheless, the analysis of PC estimators, particularly in high-dimensional linear regression, been an active area of recent research. Below, we briefly summarize some relevant contributions to provide context.
D.1 Proportional Asymptotics
Several studies analyze Principal Component Regression (PCR) in the proportional asymptotic setting where the dimension $p$ and sample size $n$ grow with $p/n\!\to\!\gamma$. In this regime, Xu and Hsu [46] study the limiting risk of PCR with Gaussian designs with diagonal covariance. They assume polynomially decaying eigenvalues or a convergent empirical spectum, together with an isotropic prior, and they reveal a “double-descent” risk curve. In a related vein, Wu and Xu [45] extend the analysis to a general covariance matrix $\Sigma_{x}$ and an anisotropic prior satisfying $\mathbb{E}\beta_{*}\beta_{*}^{\!\top}\!=\!\Sigma_{\beta}$. They also derive an exact risk expression and demonstrate how “misalignment” between $\Sigma_{x}$ and $\Sigma_{\beta}$ affects risk; here “alignment” refers to concordance between the orderings of their eigenvalues. Both studies assume knowledge of the eigenvectors of the population covariance matrix $\Sigma_{x}$ to construct the oracle PCR. Gedon et al. [17] analyze the limiting risk of PCR under a latent factor model and explore the effect of distribution shift. Green and Romanov [20] derive the exact limits of estimation risk, in-sample prediction risk, and out-of-sample prediction risk of PCR under the assumption that both the empirical spectrum distribution and the distribution of the true signal’s mass over the eigenspaces of $\Sigma_{x}$ converges weakly.
@@ D.2 Non-Asymptotic Analysis
Complementary research develops non-asymptotic guarantees. Agarwal et al. [1] derive finite-sample upper bounds on prediction error using $\|\beta^{*}\|_{1}^{2}$ and the rank of the design matrix under latent factor models, and they explore the robustness of PCR to noise and missing values in the observed covariates. Bing et al. [9] consider PCR with an adaptively selected number of components under latent factor models and provide alternative finite-sample risk bounds using $\|\beta^{*}\|_{2}^{2}$. Huang et al. [21] derive non-asymptotic risk bounds for PCR in more general settings by analyzing the alignment between population and empirical principal components. Hucker and Wahl [22] derive non-asymptotic error bounds for PCR in kernel regression.
Large Language Models Statement
Large language models were used to improve the clarity of the manuscript and to facilitate experiments.